OpenStudy (anonymous):

How do I integrate this? Please help. 2u/(1+u)

OpenStudy (anonymous):

\[\frac{ 2u }{ 1+u } du\]

OpenStudy (anonymous):

I have tried the integration by parts formula but can't get the answer. -_-

OpenStudy (anonymous):

Naw that won't work, you'll have to split it.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Never mind, just do long division, you'll get this. \[2\int\limits \left( 1 -\frac{ 1 }{ u+1 } \right)du\]

OpenStudy (anonymous):

Yeah split it now, after you do long division.

OpenStudy (anonymous):

\[2 \int\limits 1 du - 2 \int\limits \frac{ 1 }{ u+1 }du \]

OpenStudy (anonymous):

x = u+1 dx = du

OpenStudy (anonymous):

So I'll get \[2(u-\ln (1+u))\]?

OpenStudy (anonymous):

+C but yup :)

OpenStudy (anonymous):

Thank you so much. >_<

OpenStudy (anonymous):

Np :P